The table below gives the number of hours spent unsupervised each day as well as the overall grade averages for seven randomly selected middle school students. Using this data, consider the equation of the regression line, yˆ=b0+b1xy^=b0+b1x, for predicting the overall grade average for a middle school student based on the number of hours spent unsupervised each day. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Hours Unsupervised 0 0.5 1.5 2.5 4 4.5 6
Overall Grades 88 79 76 72 65 63 62

Table

Step 1 of 6:

Find the estimated slope. Round your answer to three decimal places.

Step 2 of 6:

Find the estimated y-intercept. Round your answer to three decimal places.

Step 5 of 6:

Determine if the statement "All points predicted by the linear model fall on the same line" is true or false.

Step 6 of 6:

Find the value of the coefficient of determination. Round your answer to three decimal places.
Answer #1

The equation of the regression line is

hat y = eta _{0}+eta _{1}x

where 0- Intercept and pı - sl ope , x - Hours unsupervised, y - overall grades

The formula to find the slope is,

eta _{1} =rac{nsum xy-sum xsum y}{nsum x^{2}-(sum x)^{2}}

ry- 1249 281 and n-7

By plugging the values in the formula of slope,

ηΣ.ry _ Σ.Τ Σ y 7*1249-19 * 505 7 * 81-192-=-4.136

Estimated slope = -4.136

The formula of intercept is,

β1 * (-4.136) * 83.369

Estimated intercept = 83.369

The statement is false, because the predictions will fall on the same line but the actual values will may generally not.

Coefficient of determination

First find the coefficient of correlation

The formula is,

r = rac{nsum xy-sum xsum y}{sqrt{[nsum x^{2}-(sum x)^{2}][nsum y^{2}-(sum y)^{2}]}}

a-19,y505,y 1249,2- 81,y2 36983 andn-

71249 - 19 *505 781-19-117*36983-505 3- 50 =-0.95596

Coef ficient of determination = r2 = r* r = 0.914

Coefficient of determination = 0.914